OpenStudy (anonymous):

Pages read(r) pages left to read(p) 10 65 30 45 50 25 70 5 what is the input-output function rule?

OpenStudy (anonymous):

if x = pages read and y = pages left the equation would be y = 75-x in function notation \[f(x)=75-x\]